# Feature Space Analysis by Guided Diffusion Model

### v1

@@ IV-A Results Targeting CLIP’s Image Encoder
As seen from Figs. 3 and 4, our decoder successfully generates images that have not only very similar features to $\boldsymbol{f}_{s}$ but also similar visual appearances to the actual image from which $\boldsymbol{f}_{s}$ is extracted. Quantitatively, the average pairwise squared Euclidean distance among the $15$ actual images in Figs. 3 and 4 is $6.76881$ and the one among $10$ images that are carefully chosen from the tusker category of ImageNet dataset and confirmed to have similar visual perception by human is $1.40302$ (Please see Appendix A for these $10$ images). In comparison, the squared Euclidean distance between a best image and its corresponding actual image ranges from $0.02269$ (run 2 in the bottom row of Fig. 3) to $0.14704$ (run 1 in the second row from the bottom of Fig. 4). This validates our decoder’s ability to generate images each of which has an exceptionally similar feature to $\boldsymbol{f}_{s}$.
Note that our objective is not to generate realistic images, but rather images whose features closely match $\boldsymbol{f}_{s}$. This enables us to analyse the feature space of the target feature extractor, more specifically, to investigate what images are located in proximity to the actual image from which $\boldsymbol{f}_{s}$ is extracted. Taking the first row in Fig 3 as an example, each of the generated images—like the actual image—appears to show the same type of dog with a green object around its neck, sitting outdoors with trees in the background (please enlarge the image to see details). But, the dog’s legs and tail are very unnatural, suggesting that CLIP’s image encoder pays little attention to their anatomical topology. Also, the second row in Fig. 3 implies that the feature extracted from the actual image by CLIP’s image encoder only captures a woman with red hair in a visually complex background. Please check the other rows in the same logic. This way, our decoder can offer insights in how image features are extracted by a target feature extractor (CLIP’s image encoder in this experiment) and we believe that these insights are valuable for improving it.
@@
IV-B Comparative Results Targeting ResNet-50
The main purpose of this experiment is to compare our decoder to Representation Conditional Diffusion Model (RCDM) [15]. RCDM is trained by conditioning its image generation on $\boldsymbol{f}_{s}$ so that a generated image is similar to the actual image from which $\boldsymbol{f}_{s}$ is extracted. In contrast, our decoder requires no training, instead guides the image generation of a pre-trained diffusion model (stable diffusion in this paper). Thus, our aim is to examine which of RCDM or our decoder can generate an image whose feature matches $\boldsymbol{f}_{s}$ more closely.
@@ IV-B Comparative Results Targeting ResNet-50
Figs. 5 and 6 demonstrate that our decoder generates images whose features are much closer to $\boldsymbol{f}_{s}$ than RCDM. Specifically, most of squared Euclidean distances for images generated by our decoder are single-digit while those for RCDM are three-digit. Furthermore, with respect to ResNet-50 features, the average pairwise squared Euclidean distance among the $15$ actual images in Figs. 5 and 6 is $715.830$ and the one among the $10$ tusker images carefully chosen from ImageNet dataset is $171.156$ (Please see Appendix A for these images). Based on this, when we make a rough criterion that two images whose squared Euclidean distance is less than $200$ are perceptually similar, $22$ of the $45$ images generated by RCDM in Figs. 5 and 6 are regarded as perceptually similar to their corresponding actual images. On the other hand, all the $45$ images generated by our decoder significantly surpass this criterion. Actually, although RCDM can generate images characterised by similar object layouts to the corresponding actual images, the objects in these images are not the same type. This is illustrated by, for instance, dogs in first row of Fig. 5, birds in the second row of Fig. 6 and foods in the third row from the bottom of Fig. 6.
@@
IV-C Results Targeting ViT-H14/L
Fig. 7 shows images generated by our decoder when using ViT as the target feature extractor. In particular, ViT-H/14 configured with ViT_H_14_ Weights.IMAGENET1K_SWAG_E2E_V1  [^7] https://docs.pytorch.org/vision/main/models/generated/torchvision.models.vit_h_14.html#torchvision.models.ViT_H_14_Weights is used because it is ranked at the top in the ranking of the pre-trained models in Pytorch  [^8] https://docs.pytorch.org/vision/main/models.html#table-of-all-available-classification-weights (accessed on July 30, 2025). In this ranking, ResNet-50 used in the previous section achieves a top-1 accuracy of 77.37477.374% on ImageNet-1K classification whereas ViT-H/14 reaches 88.55288.552%. The experiment below aims to examine whether this performance gap is related to images that our decoder generates by targeting ResNet-50 and ViT-H/14. Similar to the experiments before, in Fig. 7, $\boldsymbol{f}_{s}$ is extracted by ViT-H/14 from each actual image and the best images obtained by running our decoder three times are displayed next to it. In terms of ViT-H/14 features, the average pairwise squared Euclidean distance among the $15$ actual images and the one among the $10$ carefully-selected tusker images from ImageNet dataset (please see Appendix A for these images) are $2914.379$ and $715.504$, respectively. By taking these values as a reference, it can be said that our decoder successfully generates images whose features are remarkably similar to $\boldsymbol{f}_{s}$.
@@ IV-C Results Targeting ViT-H14/L
By comparing images generated by targeting ResNet-50 in Figs. 5 and 6 to the ones generated by targeting ViT-H/14 in Figs. 7, the latter clearly preserve more detailed contents in the corresponding actual images than the former. For example, the top row on the left side of Fig. 7 demonstrates that the dogs in the generated images have objects around their necks (although one of them is red), whereas such objects are absent in the generated images in the top row of Fig. 5. In addition, the second row from the bottom on the left side of Fig. 7 shows that the pizzas in the generated images and the actual one look the same. Moreover, as seen from the bottom row on the right side of Fig. 7, both of the generated images and the actual one contain tall buildings in the background (although they appear to be drawn in the sky in the left generated image). Thus, the preservation of detailed contents in images generated by our decoder can be thought to indicate the usefulness of features extracted by the target feature extractor. This corresponds to [13]’s conclusion that features (logit vectors) extracted by a better image classification model retains richer information.
@@
Regarding the comparison between CLIP’s image encoder with ResNet-50 backbone in Section IV-A and ViT-H/14 in this section, the latter is expected to outperform the former on ImageNet-1K classification. The reason is that even if a more advanced CLIP’s image encoder with ViT-Large/14 backbone is carefully finetuned on ImageNet-1K dataset, its top-1 accuracy is 88.088.0[37], which is lower than the one (88.55288.552%) of ViT-H/14 in this section. Correspondingly, the generated images in Fig. 7 seem better than the ones in Figs. 3 and 4 in terms of the preservation of detailed contents, as seen by comparing birds and foods in these images. However, the main advantage of CLIP’s image encoder is its zero-shot performance, that is, the generality of features extracted by it. As represented by the six actual images from the bottom on the right side of Fig. 7, most of images in ImageNet-1K dataset prominently display the main object near the centre. Thus, when targeting ViT-H/14 trained on such images, generated images tend to spend a large region for the main object and lose its surrounding context. In contrast, as can be seen from Figs. 3 and 4, such a surrounding context is better preserved in the images generated by targeting CLIP’s image encoder. This context preservation could be one indicator of the generality of features extracted by the target feature extractor.
IV-D Discussions
@@
This section presents four issues observed from the experiments before and discuss research topics that warrant further investigation. In particular, the first two issues are for the improvement of our decoder and the remaining two are for the construction of a better feature space.
IV-D1 Enhancing the loss function of our decoder
@@
It can be hypothesised that if we could improve our decoder to generate more realistic and natural images, those images would be located in closer proximity to the corresponding actual image in the feature space, compared images generated in this paper. As observed in images generated by targeting ViT-H/14 in Fig. 7, one typical artifact making generated images unrealistic and unnatural is large black regions, resulting from excessive modification of the predicted source noise $\epsilon_{\theta}\left(\boldsymbol{z}_{t},t\right)$ in Eq. (3) (Appendix D presents how gradient clipping mitigates this artifact). When a large region in an image is spent for such artifacts, only the remaining region can be used to make its feature match a user-specified one, which significantly limits the possibility of images that can be generated. One promising solution is to enhance the loss function by adopting a perceptual quality assessment of a generated image [38, 39] in addition to the squared Euclidean distance between its feature and the user-specified one. This reduces artifact creation and enables our decoder to use the entire image region, increasing the possibility of a close feature match.
IV-D2 Recuding the computation time of our decoder
@@
One drawback of our decoder is its computation time. Specifically, NVIDIA GeForce RTX 3090 was used to generate the images in Figs. 3 and 4 by targeting CLIP’s image encoder with ResNet-50 backbone and the average computation time per image generation run is $8830$ seconds. Using the same GPU, the images in Figs. 5 and 6 were generated by targeting ResNet-50 with an average computation time of $8951$ seconds per run. The images generated by targeting ViT-H/14 in Fig. 7 were generated using NVIDIA GeForce RTX 3090 and NVIDIA RTX 4000 Ada, which required on average $16897$ and $22694$ seconds per run, respectively. One solution for reducing these computation times  [^9] Our decoder consumes about $15.1$, $14.9$ and 19.619.6GB of VRAM for image generation runs targeting CLIP’s image encoder with ResNet-50 backbone, ResNet-50 and ViT-H/14, respectively. is to implement our decoder on a fast sampler that can be incorporated into a pre-trained diffusion model to generate images with a small number of steps [40, 41].
IV-D3 Smooth alignment between features and visual appearances
@@
IV-D4 Suitability of CLIP’s multimodal feature space
@@
Figure 8: Images generated by our decoder when using CLIP’s image encoder with ResNet-50 backbone as the target feature extractor. For the second and third columns, $\boldsymbol{f}_{s}$ is defined as the feature of each caption and its normalised one, respectively. On the other hand, $\boldsymbol{f}_{s}$ is set to $1.2$ and $0.8$ times the feature of each actual image. The number under each generated image indicates the squared Euclidean distance between its feature and $\boldsymbol{f}_{s}$ used to generated it.
Let us investigate the reason for the aforementioned result. First, CLIP is trained so that an image and the caption assigned to it are close in the feature space with respect to their (normalised) cosine similarity, not their squared Euclidean distance [6]. Thus, despite their directional similarity, they are located in completely different regions in the feature space, as verified by the fact that the average norm of the features of the six actual images in Fig. 8 and the one of the corresponding six captions are $2.177$ and $9.795$, respectively. This motivates us to run our decoder to generate an image by defining $\boldsymbol{f}_{s}$ as the caption’s normalised feature whose norm is equal to the average feature norm of the six actual images. Nonetheless, the column named “Normalised” in Fig. 8 shows that the generated images significantly differ from the corresponding actual images, although they slightly align with the captions as typically observed in the second image from the bottom.
@@
V Conclusion and Future Work
This paper introduced a decoder that generates images whose features are ensured to closely match user-specified ones. Owing to the short Euclidean distances between these features, we can perform rigorous analysis of the feature space of a target feature extractor. Our decoder is implemented as a guided diffusion model that needs no training, instead guides a pre-trained diffusion model (stable diffusion in this paper) to minimise the squared Euclidean distance between the feature of a generated image and the user-specified one. Furthermore, the early step emphasis strategy and gradient normalisation and clipping are devised to improve our decoder’s image generation. The experiments using each of CLIP’s image encoder with ResNet-50 backbone, ResNet-50 and ViT-H/14 as the target feature extractor demonstrate that images generated by our decoder have features remarkably similar to the specified ones and offer valuable insights into the feature space of the extractor.
In addition to the four issues discussed in Section IV-D, we plan to implement our decoder on pre-trained diffusion models other than stable diffusion and to analyse feature spaces of more recent feature extractors than the ones employed in this paper (e.g., image encoders implemented in OpenCLIP project [42]). Also, we aim to improve our decoder’s image generation by formulating the guidance-based reverse process as a search problem to find the optimal sequence of denoising operations [43, 44]. Finally, our encoder can be readily extended to other modalities like text, audio and time-series to analyse their respective feature spaces.
Appendix A $10$ Carefully-chosen Tusker Images

### v2

@@ 4.1 Results Targeting CLIP’s Image Encoder
As seen from Figs. 3 and 4, our decoder successfully generates images that have not only very similar features to $\boldsymbol{f}_{s}$ but also similar visual appearances to the actual image from which $\boldsymbol{f}_{s}$ is extracted. Quantitatively, the average pairwise squared Euclidean distance among the $15$ actual images in Figs. 3 and 4 is $6.76881$ and the one among $10$ images that are carefully chosen from the tusker category of ImageNet dataset and confirmed to have similar visual perception by human is $1.40302$ (Please see Part A in the supplementary materials for these $10$ images). In comparison, the squared Euclidean distance between a best image and its corresponding actual image ranges from $0.02269$ (run 2 in the bottom row of Fig. 3) to $0.14704$ (run 1 in the second row from the bottom of Fig. 4). This validates our decoder’s ability to generate images each of which has an exceptionally similar feature to $\boldsymbol{f}_{s}$.
Note that our objective is not to generate realistic images, but rather images whose features closely match $\boldsymbol{f}_{s}$. This enables us to analyse the feature space of the target feature extractor, more specifically, to investigate what images are located in proximity to the actual image from which $\boldsymbol{f}_{s}$ is extracted. Taking the first row in Fig 3 as an example, each of the generated images—like the actual image—appears to show the same type of dog with a green object around its neck, sitting outdoors with trees in the background (please enlarge the image to see details). But, the dog’s legs and tail are very unnatural, suggesting that CLIP’s image encoder pays little attention to their anatomical topology. Also, the second row in Fig. 3 implies that the feature extracted from the actual image by CLIP’s image encoder only captures a woman with red hair in a visually complex background. Please check the other rows in the same logic. This way, our decoder can offer insights into how image features are extracted by a target feature extractor (CLIP’s image encoder in this experiment) and we believe that these insights are valuable for improving it.
@@
4.2 Comparative Results Targeting ResNet-50
The main purpose of this experiment is to compare our decoder to Representation Conditional Diffusion Model (RCDM) [15]. RCDM is trained by conditioning its image generation on $\boldsymbol{f}_{s}$ so that a generated image is similar to the actual image from which $\boldsymbol{f}_{s}$ is extracted. In contrast, our decoder requires no training, instead guides the image generation of a pre-trained diffusion model (stable diffusion in this paper). Our aim is to examine which of RCDM or our decoder can generate an image whose feature matches $\boldsymbol{f}_{s}$ more closely.
@@ 4.2 Comparative Results Targeting ResNet-50
Figs. 5 and 6 demonstrate that our decoder generates images whose features are much closer to $\boldsymbol{f}_{s}$ than RCDM. Specifically, most of squared Euclidean distances for images generated by our decoder are single-digit while those for RCDM are three-digit. With respect to ResNet-50 features, the average pairwise squared Euclidean distance among the $15$ actual images in Figs. 5 and 6 is $715.830$ and the one among the $10$ tusker images carefully chosen from ImageNet dataset is $171.156$ (Please see Part A in the supplementary materials for these images). Based on this, when we make a rough criterion that two images whose squared Euclidean distance is less than $200$ are perceptually similar, $22$ of the $45$ images generated by RCDM in Figs. 5 and 6 are regarded as perceptually similar to their corresponding actual images. On the other hand, all the $45$ images generated by our decoder significantly surpass this criterion. Actually, although RCDM can generate images characterised by similar object layouts to the corresponding actual images, the objects in these images are not the same type. This is illustrated by, for instance, dogs in first row of Fig. 5, birds in the second row of Fig. 6 and foods in the third row from the bottom of Fig. 6.
@@
4.3 Results Targeting ViT-H14/L
Fig. 7 shows images generated by our decoder when using ViT as the target feature extractor. In particular, ViT-H/14 configured with ViT_H_14_ Weights.IMAGENET1K_SWAG_E2E_V1  [^7] https://docs.pytorch.org/vision/main/models/generated/torchvision.models.vit_h_14.html#torchvision.models.ViT_H_14_Weights is used because it is ranked at the top in the ranking of the pre-trained models in Pytorch  [^8] https://docs.pytorch.org/vision/main/models.html#table-of-all-available-classification-weights (accessed on July 30, 2025). In this ranking, ResNet-50 used in the previous section achieves a top-1 accuracy of 77.37477.374% on ImageNet-1K classification whereas ViT-H/14 reaches 88.55288.552%. The experiment below aims to examine whether this performance gap is related to images that our decoder generates by targeting ResNet-50 and ViT-H/14. Similar to the experiments before, in Fig. 7, $\boldsymbol{f}_{s}$ is extracted from each actual image and the best images obtained by running our decoder three times are displayed. In terms of ViT-H/14 features, the average pairwise squared Euclidean distance among the $15$ actual images and the one among the $10$ carefully-selected tusker images from ImageNet dataset (please see Part A in the supplementary materials for these images) are $2914.379$ and $715.504$, respectively. By taking these values as a reference, it can be said that our decoder successfully generates images whose features are remarkably similar to $\boldsymbol{f}_{s}$.
@@ 4.3 Results Targeting ViT-H14/L
By comparing images generated by targeting ResNet-50 in Figs. 5 and 6 to the ones generated by targeting ViT-H/14 in Fig. 7, the latter clearly preserve more detailed contents in the corresponding actual images than the former. For example, the top row on the left side of Fig. 7 demonstrates that the dogs in the generated images have objects around their necks (although one of them is red), whereas such objects are absent in the generated images in the top row of Fig. 5. In addition, the second row from the bottom on the left side of Fig. 7 shows that the pizzas in the generated images and the actual one look the same. Moreover, as seen from the bottom row on the right side of Fig. 7, both of the generated images and the actual one contain tall buildings in the background (although they appear to be drawn in the sky in the left generated image). Thus, the preservation of detailed contents in images generated by our decoder can be thought to indicate the usefulness of features extracted by the target feature extractor. This corresponds to [13]’s conclusion that features (logit vectors) extracted by a better image classification model retains richer information.
@@
Regarding the comparison between CLIP’s image encoder with ResNet-50 backbone in Section 4.1 and ViT-H/14 in this section, the latter is expected to outperform the former on ImageNet-1K classification. The reason is that even if a more advanced CLIP’s image encoder with ViT-Large/14 backbone is carefully finetuned on ImageNet-1K dataset, its top-1 accuracy is 88.088.0[37], which is lower than the one (88.55288.552%) of ViT-H/14 in this section. Correspondingly, the generated images in Fig. 7 seem better than the ones in Figs. 3 and 4 in terms of the preservation of detailed contents, as seen by comparing birds and foods in these images. However, the main advantage of CLIP’s image encoder is its zero-shot performance, that is, the generality of features extracted by it. As represented by the six actual images from the bottom on the right side of Fig. 7, most of images in ImageNet-1K dataset prominently display the main object near the centre. Thus, when targeting ViT-H/14 trained on such images, generated images tend to spend a large region for the main object and lose its surrounding context. In contrast, as can be seen from Figs. 3 and 4, such a surrounding context is better preserved in the images generated by targeting CLIP’s image encoder. This context preservation could be one indicator of the generality of features extracted by the target feature extractor.
4.4 Discussions
@@
This section presents four issues observed from the experiments before and discuss research topics that warrant further investigation. The first two issues are for the improvement of our decoder and the remaining two are for the construction of a better feature space.
4.4.1 Enhancing the loss function of our decoder
@@
It can be hypothesised that if we could improve our decoder to generate more realistic and natural images, those images would be located in closer proximity to the corresponding actual image in the feature space, compared images generated in this paper. As observed in images generated by targeting ViT-H/14 in Fig. 7, one typical artifact making generated images unrealistic and unnatural is large black regions, resulting from excessive modification of the predicted source noise $\epsilon_{\theta}\left(\boldsymbol{z}_{t},t\right)$ in Eq. (3) (Part D in the supplementary materials presents how gradient clipping mitigates this artifact). When a large region in an image is spent for such artifacts, only the remaining region can be used to make its feature match a user-specified one, which significantly limits the possibility of images that can be generated. One promising solution is to enhance a loss function by adopting a perceptual quality assessment of a generated image [38, 39] in addition to the squared Euclidean distance between its feature and the user-specified one. This reduces artifact creation and enables our decoder to use the entire image region, thereby increasing the possibility of a close feature match.
4.4.2 Recuding the computation time of our decoder
@@
One drawback of our decoder is its computation time. Specifically, NVIDIA GeForce RTX 3090 was used to generate the images in Figs. 3 and 4 by targeting CLIP’s image encoder with ResNet-50 backbone and the average computation time per image generation run is $8830$ seconds. Using the same GPU, the images in Figs. 5 and 6 were generated by targeting ResNet-50 with an average computation time of $8951$ seconds per run. The images generated by targeting ViT-H/14 in Fig. 7 were generated using NVIDIA GeForce RTX 3090 and NVIDIA RTX 4000 Ada, which required on average $16897$ and $22694$ seconds per run, respectively. One solution for reducing our decoder’s computation time  [^9] Our decoder consumes about $15.1$, $14.9$ and 19.619.6GB of VRAM for image generation runs targeting CLIP’s image encoder with ResNet-50 backbone, ResNet-50 and ViT-H/14, respectively. is to adopt a fast sampler that can be incorporated into a pre-trained diffusion model to generate images with a small number of steps [40, 41].
4.4.3 Smooth alignment between features and visual appearances
@@
4.4.4 Suitability of CLIP’s multimodal feature space
@@
Figure 8: Images generated by our decoder when using CLIP’s image encoder with ResNet-50 backbone as the target feature extractor. For the second and third columns, $\boldsymbol{f}_{s}$ is defined as the feature of each caption and its normalised one, respectively. On the other hand, for the forth and last columns, $\boldsymbol{f}_{s}$ is set to $1.2$ and $0.8$ times the feature of each actual image, respectively. The number under each generated image and the number in brackets (if present) indicate the squared Euclidean distance and the normalised cosine similarity between its feature and $\boldsymbol{f}_{s}$ used to generated it, respectively.
Let us investigate the reason for the aforementioned result. CLIP is trained so that an image and the caption assigned to it are close in the feature space with respect to their (normalised) cosine similarity, not their squared Euclidean distance [6]. For the six actual image-caption pairs in Fig. 8, the average cosine similarity between their features is $0.224$ (the maximum $0.266$). In contrast, the cosine similarity between the feature of each generated image and the one of the corresponding caption is significantly higher, as indicated by the numbers in brackets in Fig. 8  [^10] When defining $\boldsymbol{f}_{s}$ as the feature of an actual image, the cosine similarity between $\boldsymbol{f}_{s}$ and the feature of a generated image is always larger than $0.99$, so such cosine similarities are omitted in Fig. 8.. This means that, in the feature space, the generated image is oriented much closer to the caption than the paired actual image. Nonetheless, they are located in completely different regions as verified by the fact that the average norm of the features of the six actual images in Fig. 8 and the one of the corresponding six captions are $2.177$ and $9.795$, respectively. This motivates us to run our decoder to generate an image by defining $\boldsymbol{f}_{s}$ as the caption’s normalised feature whose norm is equal to the average feature norm of the six actual images. However, the column named “Normalised” in Fig. 8 shows that the generated images significantly differ from the corresponding actual images, although they slightly align with the captions as typically observed in the second image from the bottom.
@@
5 Conclusion and Future Work
This paper introduced a decoder that generates images whose features are ensured to closely match user-specified ones. Owing to the short Euclidean distances between these features, we can perform rigorous analysis of the feature space of a target feature extractor. Our decoder is implemented as a guided diffusion model that needs no training, instead guides a pre-trained diffusion model (stable diffusion in this paper) to minimise the squared Euclidean distance between the feature of a generated image and the user-specified one. Furthermore, the early step emphasis strategy and gradient normalisation and clipping are devised to improve our decoder’s image generation. The experiments using each of CLIP’s image encoder with ResNet-50 backbone, ResNet-50 and ViT-H/14 as the target feature extractor demonstrate that images generated by our decoder have features remarkably similar to the user-specified ones and offer valuable insights into the feature space of the extractor.
In addition to the four issues discussed in Section 4.4, we plan to implement our decoder on pre-trained diffusion models other than stable diffusion and to analyse feature spaces of more recent feature extractors than the ones employed in this paper (e.g., image encoders implemented in OpenCLIP project [42]). Also, we aim to improve our decoder’s image generation by formulating the guidance-based reverse process as a search problem to find the optimal sequence of denoising operations [43, 44]. Finally, our encoder can be readily extended to other modalities like text, audio and time-series to analyse their respective feature spaces.
6 Declaration of Generative AI and AI-assisted Technologies in the Manuscript Preparation Process
During the preparation of this work, the author(s) used ChatGPT (GPT-5) only to revise sentences that had been originally written by the authors. In other words, no sentence in this paper was written solely by ChatGPT. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Supplementary Materials
Part A $10$ Carefully-chosen Tusker Images